How do I find the surface area if my room measurements are as follows
Length 8ft 5"
Width 5ft 3"
Height 10ft
I understand the formula is:
SA=2lw+2lh+2wh
Am I supposed to convert these to inches first and then solve with the formula? If not, please let me know how I would multiply 8ft 5inches x 10ft - how would that come out correctly?

Found 2 solutions by unlockmath, rapaljer:
Answer by unlockmath(1688)

Hello,
I'll give you a hint and then you can plug in the numbers. Change the ft & inches to a fraction: for example 8 ft 5 in would be 8 and 5/12 ft. When you've done that then you can multiply.
Make sense?
RJ

Answer by rapaljer(4671)

I recommend that you convert the inches to feet, and work the problem in decimal format. You have a calculator don't you?
5 inches = 5/12 = .416666 feet
3 inches = 3/12 = .25 feet

Multiply 8.41666666*5.25, etc.

Dr. Robert J. Rapalje, Retired